OpenStudy (anonymous):

determine whether the give point is on the circle:: P(-2,3) on 2x^2+2y^2=26

mathslover (mathslover):

Can you tell me, what will be the radius of the circle?

OpenStudy (anonymous):

2sqrt2?

mathslover (mathslover):

can you tell me how did you do that?

OpenStudy (anonymous):

i used the distance formula

mathslover (mathslover):

do you know, what should be the condition for a point to be inside a circle?

mathslover (mathslover):

* do you know, what should be the condition for a point to be on a circle?

mathslover (mathslover):

See first of all we are given with the equation of a circle as : 2x^2 + 2y^2 = 26 or : 2(x^2 + y^2) = 26 x^2 + y^2 = 13 or x^2 + y^2 = \(\sqrt{13}^2\) Hence, 13 is considered as the radius of the circle with origin as the centre.

OpenStudy (anonymous):

ohhhhhhhhh okok

OpenStudy (anonymous):

should i still place the sqrt in 13? or no......

mathslover (mathslover):

Though, you don't need to do that for solving this question. We have : \(x^2 + y^2 = 13\\ (-2)^2 + (3)^2 = 13\\ 4 + 9 = 13\\ 13=13 \) Which is true^ , thus the point lies ON the circle.

mathslover (mathslover):

The condition is : x^2 + y^2 = (radius)^2 Taking centre as origin then for (x,y) point to lie "ON" the circle , x^2 + y^2 must be equal to (radius)^2

OpenStudy (anonymous):

kamsahamnida!!!!!!

mathslover (mathslover):

Sorry, what does that mean? ^

OpenStudy (anonymous):

omg sry it means thank u :)

mathslover (mathslover):

:D You're welcome. Good Luck

Directrix (directrix):

감사합니다